Kevin drew a regular polygon the measure of one of its interior angles explain why is impossible for his measure to be 145 degrees
Answer by Alan3354(69443)

Kevin drew a regular polygon the measure of one of its interior angles explain why is impossible for his measure to be 145 degrees
==================
Ext angles = 180 - 145 = 35 degs
# of sides of ALL polygons = 360/Ext angle
360/35 is not an integer
--> not possible